If I had 25g of Propane, how many grams of Carbon and Hydrogen would I have?
1
Expert's answer
2020-04-18T07:06:50-0400

If I had 25g of Propane, how many grams of Carbon and Hydrogen would I have?

M(C3H8) = 44 g/mol

n(C3H8) = m/M

n(C3H8) = 25/44 = 0.57 mol

n(C) = 3 * 0.57 = 1.7 mol

n(H) = 8 * 0.57 = 4.6 mol

m = n*M

m(C) = 1.7 * 12 = 20.4g.

m (H) = 4.6 * 1 = 4.6g.

Answer: m(C) = 20.4g, m(H) = 4.6g.